# Do global forecasting models require frequent retraining?

### v1

@@ 1 Introduction
Forecasting plays a critical role in decision-making processes across industries, from supply chain management to energy demand planning. Traditionally, time series forecasting has relied on a Local Modeling approach (LM), meaning that separate models 111 We use the words model and method interchangeably even if the two have distinct statistical meanings (Svetunkov & Boylan, 2023). are trained for each series in isolation. However, recent advances have demonstrated the potential of Global forecasting Models (GM), a paradigm also known as cross-learning, which in turn consists of fitting a single forecasting model on the whole data set (Januschowski et al., 2020). This novel approach allows to leverage information across multiple time series to improve accuracy and generalization by employing machine learning or deep learning architectures 222 We distinguish between traditional machine learning models, like linear regression or tree-based methods, and deep learning models, those that are based on some sort of neural network architecture. . Global modeling has affirmed itself as one of the most relevant innovations in time series forecasting of recent years (Semenoglou et al., 2021), and its success is not only due to the demonstrated forecast accuracy in most time series competitions, but also to the particular computational benefit it provides. Indeed, while the local approach to time series forecasting is potentially beneficial for forecast accuracy, it comes with substantial computational overhead. Moreover, since the forecasts are usually produced through some cloud computing service that relies on pay-as-you-go pricing (Fotios Petropoulos & Spiliotis, 2024), higher computing time and resources directly translate into higher costs of forecasting for organizations.
@@
Nevertheless, despite the growing adoption of global models, it is still common practice to update the forecasting models when new observations are available, often motivated by the assumption that continuous updates lead to better adaptability to changing patterns and more accurate predictions. Unlike local models, global models benefit from learning shared dynamics across multiple time series, potentially leading to more robust representations that are less sensitive to frequent retraining. In the context of local models, Spiliotis & Petropoulos (2024) demonstrated that less frequent model updates do not harm forecasting accuracy. However, in the context of global models, the effects of retraining are not well understood. Is continuous retraining necessary to maintain forecast performance, or can global models remain effective without frequent updates? This aspect remains largely unexplored in the forecasting literature and investigating the impact of retraining is crucial for both methodological and practical reasons. From a methodological perspective, understanding whether global models degrade in performance without frequent retraining provides insights into their stability and adaptability. If global models can maintain strong predictive accuracy with less frequent updates, it would challenge conventional wisdom on the necessity of continuous retraining in forecasting. Moreover, from a practical perspective, frequent retraining has significant computational and environmental costs. Training large-scale forecasting models requires substantial computational power, contributing to energy consumption and carbon emissions (Schwartz et al., 2020). Indeed, the energy consumption of model training extends beyond direct computational costs. In recent years, the environmental impact of machine learning models has become a growing concern. Data centers running these models contribute significantly to global carbon emissions, with estimates suggesting that training a single large deep learning model can emit as much carbon as five cars over their lifetimes Strubell et al. (2019). The frequency of model retraining multiplies this impact, particularly in large organizations handling millions of series, emphasizing the need for more computationally efficient training practices. The implication is straightforward: reducing the retraining frequency of forecasting models could contribute substantially to sustainability by lowering energy consumption, possibly without harming accuracy performance.
@@ 1.1 Research Question
We aim to address the question ”Is frequent retraining necessary in the context of global forecasting models?”. Specifically, we study whether skipping retraining when new observations are available harms the forecasting performance of global models. To answer this question, we use the two most recent and comprehensive retail forecasting datasets, namely the M5 and the VN1 competitions’ data.
To generally understand how retraining the forecasting models affects their performances, we consider ten different global forecasting methods (five from the ”classical” machine learning domain and five often used deep neural network architectures), and several possible retraining scenarios, from continuous retraining to no retraining at all. We also explore intermediate periodic retraining strategies to broadly cover the most reasonable and effective scenarios.
@@ 1.2 Contributions
    We compare different retraining scenarios (e.g., continuous, periodic, and no retraining) on different datasets to quantify the impact of frequent retraining in terms of cost of forecasting.
@@ 1.3 Overview
The rest of this paper is organized as follows. After a brief review of related works (Section 2), in Section 3 we describe the design of the experiment used in our study. The datasets and their characteristics are presented in 3.1, the methods adopted for global forecasting are discussed in 3.2, and the concepts related to model update and retrain scenario are explained in 3.3. The evaluation setup through rolling origin validation is presented in 3.4, and the metrics used to evaluate the model performances are shown in 3.5. In Section 4 we discuss the empirical findings of our study, including forecast accuracy, computing time, and cost analysis of the different scenarios. Finally, Section 5 contains our summary and conclusions.
@@ 2 Related works
The literature on cross-learning approach has evolved significantly in recent years. Nowadays, most of the works related to time series forecasting include at least some benchmark comparison with global models, demonstrating their relevance in the field. Semenoglou et al. (2021) extensively showed their accuracy on the M4 competition dataset, Hewamalage et al. (2022) evaluated the conditions when global forecasting models are competitive, and Montero-Manso & Hyndman (2021) and (Montero-Manso, 2023) theoretically demonstrated that GM are at least as accurate as local models with less complexity and without any assumption on the similarity of the data. GM emerged as the most accurate approach in many forecasting areas, such as gas consumption (Gaweł & Paliński, 2024), electricity demand (Buonanno et al., 2022), water demand (Groß & Hans, 2024), crop production (Ibañez & Monterola, 2023), and retail demand (Spiliotis et al. (2022), Bandara et al. (2019), Juan R Trapero & Fildes (2015)). Moreover, GM effectiveness shined during the M5 competition (Makridakis et al., 2022a), where tree-based models leveraging cross-learning were among almost all the top forecasting solutions (Januschowski et al., 2022). Several techniques like clustering (Godahewa et al. (2021a), Bandara et al. (2020)), and data augmentation (Bandara et al., 2021) have been tested to further increase the performances of GM. Furthermore, new machine learning (Godahewa et al., 2023) and deep learning (Oreshkin et al., 2020) architectures specifically designed for cross-learning have been developed. Finally, a novel area of research is emerging that tries to improve the ability of GM to capture local patterns (Sen et al., 2019), and their explainability (Rajapaksha et al., 2023).
@@
From a forecasting evaluation perspective, most of the literature on GM is focused on point prediction accuracy, possibly because most machine learning and deep learning methods do not directly output probabilistic forecasts (Makridakis et al., 2022c). Nevertheless, in many forecasting contexts (like supply chain) it is very important to be able to produce and evaluate predictions in a probabilistic way (being intervals, quantiles or density based) (Fildes et al., 2022). Among others, Vovk et al. (2005) introduced a novel tool for uncertainty quantification under any machine learning model, namely Conformal Inference, that can also be applied in time series forecasting experiments (Stankeviciute et al., 2021).
@@
In the context of model retraining and updating strategies, instead, Spiliotis & Petropoulos (2024) is the main work related to time series forecasting. The authors explored extensively the effects of different retraining scenarios and different forms of model parameter updates on the model performance, although, they focused on the exponential smoothing family of models following the traditional local approach. (Huber & Stuckenschmidt, 2020) briefly discussed retraining in the context of retail demand, but with few models and retraining scenarios, and on a proprietary daily dataset only. Despite the findings being promising, there has been little direct exploration of whether global models specifically require frequent retraining or if they can retain competitive accuracy with less frequent updates. However, this topic is under consideration of the broader machine learning community (Getzner et al., 2023), advocating for Green AI (Schwartz et al., 2020).
@@
Our study builds upon these existing works, directly investigating the necessity of retraining in global forecasting models. By evaluating many different retraining strategies and their impact on the forecast accuracy of several global models, we aim to provide both theoretical insights and practical recommendations for sustainable forecasting practices.
@@ 3 Experimental design
This section describes the empirical analysis we conducted to study whether less frequent retraining scenarios may produce similar accuracy results concerning the baseline scenario (that with the highest retraining frequency). First, we describe the datasets used in the experiments, and then explain the different machine learning and deep learning models adopted. The performance measures, the several possible scenarios, and the strategy used to evaluate the forecasts are also discussed.
@@ 3.1 Datasets
For the experiment, we used two retail forecasting datasets: the M5 and the VN1 competition datasets. The M5 competition was organized by Spyros Makridakis and his colleagues as part of the M-competitions series, which aimed to compare different forecasting methods in the context of retail demand (Makridakis et al., 2022b). The M5 (Howard & Makridakis, 2020) is a well-known and well-studied dataset containing 3.049 daily time series of Walmart’s unit sales of products. It covers the sales of three categories of products (Food, Hobbies, and Household) sold into ten different stores located in three US states (California, Texas, and Wisconsin). The time period spans from 2011 until 2016. The time series are highly intermittent and are hierarchically organized, allowing forecasting at multiple levels such as individual products, product categories, stores, and States. Exogenous information that can influence sales, such as product prices, promotions, and special events (e.g., holidays) are also available. The VN1 Forecasting - Accuracy Challenge competition was jointly organized by Flieber, Syrup Tech, and SupChains in October 2024, and it is the first of its series (Vandeput, 2024). The dataset contains weekly sales of 15.053 products sold from 2020 until 2024 from e-vendors. In particular, they were mostly online from the US, and the products were directly sold to the final consumers. Contrary to the M5 dataset, where all the products were sold by a single retailer (Walmart) and from just a few stores, the VN1 dataset includes product sales of 328 warehouses from 46 different retailers. As far as we know, we are among the first to test forecasting models on this data. These sets of data are the most recent and comprehensive time series datasets related to retail demand, allowing for a good generalization of the results in the context of demand forecasting.
Table 1: Characteristics of the different datasets used in the experiments.
@@
In both cases, in our experiment, we focused on the most disaggregated level (SKUs), since the potential benefits of retraining the models less frequently are much larger at lower levels of aggregation. Moreover, we did not consider the whole set of time series to be able to consistently apply the evaluation setup described in Sections 3.4. In particular, for daily data, we kept only time series with at least two years of data (730 observations), while for weekly SKUs we considered only those that had at least three years of data (157 observations).
@@ 3.2 Forecasting models
Let define $\mathcal{Y}$ as the set of all available time series in a dataset, such that $Y_{i}$ represents a single component, and let $\mathcal{F}$ be the set of possible predictive functions, so that $F$ corresponds to a single model 444 In our setting, $F$ can be any predictive model in the machine learning or deep learning framework. We do not consider classical statistical forecasting methods, like ARIMA or Exponential Smoothing, since in their common formulations they are purely local models. . Without loss of generality, we can assume that all the necessary information for prediction are contained in $\mathcal{Y}$. Under the local approach, predictions for the forecast horizon $h$ are obtained training a model for each time series in the data set, implying that each time series has its own model, defined by its own parameters’ values.
@@
On the contrary, following the global modeling framework, forecasts for are each time series are produced by a model trained on the whole data set.
@@
In our experiment, we focused on analyzing the performances of global methods only since we are mainly interested in testing whether this approach, as opposed to the local one, can benefit from retraining the models less frequently. Indeed, nowadays cross-learning is the go-to approach for most industries that are involved with huge time series datasets, as in the context of retail demand forecasting, where the forecasts of many different SKUs have to be provided regularly (Januschowski et al., 2020). For a comprehensive evaluation of global modeling approaches in demand forecasting we used both traditional machine learning models and cutting-edge deep learning methodologies. The models were selected for their established performance in time series forecasting tasks and their diverse methodological approaches, enabling a wide comparison.
@@ 3.2.1 Machine learning models
Machine learning models have demonstrated their effectiveness in forecasting tasks due to their ability to capture non-linear relationships in the data. Moreover, they are often easy to train and usually produce very accurate results leveraging the cross-learning approach. In this study, we experimented with Linear (Pooled) Regression and four different tree-based methods.
Linear Regression (LR) is a classical statistical model that assumes a linear relationship between input features and the target variable. Despite its simplicity, LR can be effective for time series forecasting when combined with appropriate feature engineering. The Pooled Regression is considered a solid benchmark for global model performance evaluation ((Montero-Manso & Hyndman, 2021), (Godahewa et al., 2021b)), and it has also proven to be quite effective (Bandara et al., 2022).
Random Forest (RF) is an ensemble learning method based on regression trees. It constructs multiple trees during training and aggregates their predictions through averaging. RF excels in capturing non-linear patterns and interactions between variables and it is robust to overfitting making it a strong tree-based model for demand forecasting (Breiman, 2001). It was the method used by Amazon until 2015 to forecast e-commerce products demand (Januschowski et al., 2022) 555 For the cost of computation, the Random Forest model is tested only on the VN1 weekly dataset. .
@@
Light Gradient Boosting Machine (LGBM) is another gradient-boosted tree algorithm that emphasizes efficiency and scalability. It uses a histogram-based approach to split features and a leaf-wise growth strategy to reduce computation time. LGBM is particularly well-suited for large datasets and high-dimensional data, often outperforming other boosting algorithms in terms of speed and accuracy (Ke et al., 2017). In demand forecasting, it is among the top solutions of all the major competitions (Favorita, Rossmann, M5, and VN1) (Makridakis et al. (2022a), Makridakis et al. (2022c), In & Jung (2022), Lainder & Wolfinger (2022)).
Categorical Boosting (CatBoost) is a gradient-boosted decision tree algorithm specifically designed to handle categorical data. By leveraging ordered boosting and other innovations it fosters accuracy and enhances generalization (Prokhorenkova et al., 2018). CatBoost’s ability to handle categorical features without extensive preprocessing makes it advantageous for forecasting tasks involving categorical covariates. During the last years, it consistently showed top performance on many tabular data studies, becoming the go-to solution for many practitioners in the field (Shmuel et al. (2024), Ye et al. (2025)).
Machine learning models have the advantage of being easier to train with respect to their deep learning alternatives. However, they require extensive and careful feature engineering to produce accurate results (Januschowski et al., 2022). For this reason, we followed simplified versions of the M5 and VN1 top solutions to build time series features. In particular, we used lags, rolling averages, expanding averages, calendar features (year, quarter, month, week, day of week, day) and static features (store, category, location, product identifiers) based on the frequency of the dataset. Moreover, for the M5 data, we used also external features related to special events since available. The most relevant hyperparameters of the models were selected based on top performant solutions, otherwise the default values suggested by the software provider were adopted.

### v2

@@ 1 Introduction
Forecasting plays a critical role in decision-making processes across industries, from supply chain management to energy demand planning. Traditionally, time series forecasting has relied on a Local Modeling approach (LM), meaning that separate models 111 We use the words model and method interchangeably, even if the two have distinct statistical meanings (Svetunkov & Boylan, 2023). are trained for each series in isolation. However, recent advances have demonstrated the potential of Global forecasting Models (GM), a paradigm also known as cross-learning, which in turn consists of fitting a single forecasting model on the whole data set (Januschowski et al., 2020). This novel approach allows to leverage information across multiple time series to improve accuracy and generalization by employing machine learning or deep learning architectures 222 We distinguish between traditional machine learning models, like linear regression or tree-based methods, and deep learning models, those that are based on some sort of neural network architecture. . Global modeling has affirmed itself as one of the most relevant innovations in time series forecasting of recent years (Semenoglou et al., 2021), and its success is not only due to the demonstrated forecast accuracy in most time series competitions, but also to the particular computational benefit it provides. Indeed, while the local approach to time series forecasting is potentially beneficial for forecast accuracy, it comes with substantial computational overhead. Moreover, since the forecasts are usually produced through some cloud computing service that relies on pay-as-you-go pricing (Fotios Petropoulos & Spiliotis, 2024), higher computing time and resources directly translate into higher costs of forecasting for organizations.
@@
Nevertheless, despite the growing adoption of global models, it is still common practice to update the forecasting models when new observations are available, often motivated by the assumption that continuous updates lead to better adaptability to changing patterns and more accurate predictions. Unlike local models, global models benefit from learning shared dynamics across multiple time series, potentially leading to more robust representations that are less sensitive to frequent retraining. In the context of local models, the seminal work of Spiliotis & Petropoulos (2024) demonstrated that less frequent model updates do not harm forecasting accuracy. However, in the context of global models, the effects of retraining are not well understood. Is continuous retraining necessary to maintain forecast performance, or can global models remain effective without frequent updates? This aspect remains largely unexplored in the forecasting literature, and investigating the impact of retraining is crucial for both methodological and practical reasons. From a methodological perspective, understanding whether global models degrade in performance without frequent retraining provides insights into their stability and adaptability. If global models can maintain strong predictive accuracy with less frequent updates, it would challenge conventional wisdom on the necessity of continuous retraining in forecasting. Moreover, a stable forecasting model provides great benefit to the whole decision process the forecasts are used for. From a practical perspective, frequent retraining has significant computational and environmental costs. Training large-scale forecasting models requires substantial computational power, contributing to energy consumption and carbon emissions (Schwartz et al., 2020). Indeed, the energy consumption of model training extends beyond direct computational costs. In recent years, the environmental impact of machine learning models has become a growing concern. Data centers running these models contribute significantly to global carbon emissions, with estimates suggesting that training a single large deep learning model can emit as much carbon as five cars over their lifetimes (Strubell et al., 2019). The frequency of model retraining multiplies this impact, particularly in large organizations handling millions of series, emphasizing the need for more computationally efficient training practices. The implication is straightforward: reducing the retraining frequency of forecasting models could contribute substantially to sustainability by lowering energy consumption, possibly without harming accuracy performance.
@@ 1.1 Research Question
We aim to extend the results presented in the influential work of Spiliotis & Petropoulos (2024), addressing the question ”Is frequent retraining necessary in the context of global forecasting models?”. Specifically, we study whether skipping retraining when new observations are available harms the forecasting performance of global models. To answer this question, we use the two most recent and comprehensive retail forecasting datasets, namely the M5 and the VN1 competitions’ data.
To generally understand how retraining the forecasting models affects their performance, we consider ten different global forecasting methods (five from the ”classical” machine learning domain and five often-used deep neural network architectures), and several possible retraining scenarios, from continuous retraining to no retraining at all. We also explore intermediate periodic retraining strategies to broadly cover the most reasonable and effective scenarios.
@@ 1.2 Contributions
    We compare different retraining scenarios (e.g., continuous, periodic, and no retraining) on different datasets to quantify the impact of frequent retraining in terms of the cost of forecasting.
@@ 1.3 Overview
The rest of this paper is organized as follows. After a brief review of related works (Section 2), in Section 3, we describe the design of the experiment used in our study. The datasets and their characteristics are presented in 3.1, the methods adopted for global forecasting are discussed in 3.2, and the concepts related to model update and retrain scenario are explained in 3.3. The evaluation setup through rolling origin validation is presented in 3.4, and the metrics used to evaluate the model performances are shown in 3.5. In Section 4, we discuss the empirical findings of our study, including forecast accuracy, computing time, and cost analysis of the different scenarios. Finally, Section 5 contains our summary and conclusions.
@@ 2 Related works
The literature on cross-learning approach has evolved significantly in recent years. Nowadays, most of the works related to time series forecasting include at least some benchmark comparison with global models, demonstrating their relevance in the field. Semenoglou et al. (2021) extensively showed their accuracy on the M4 competition dataset, Hewamalage et al. (2022) evaluated the conditions when global forecasting models are competitive, and Montero-Manso & Hyndman (2021) and Montero-Manso (2023) theoretically demonstrated that GM are at least as accurate as local models with less complexity and without any assumption on the similarity of the data. GM emerged as the most accurate approach in many forecasting areas, such as gas consumption (Gaweł & Paliński, 2024), electricity demand (Buonanno et al., 2022), water demand (Groß & Hans, 2024), crop production (Ibañez & Monterola, 2023), and retail demand (Spiliotis et al. (2022), Bandara et al. (2019), Juan R Trapero & Fildes (2015)). Moreover, GM effectiveness shined during the M5 competition (Makridakis et al., 2022a), where tree-based models leveraging cross-learning were among almost all the top forecasting solutions (Januschowski et al., 2022). Several techniques like clustering (Godahewa et al. (2021a), Bandara et al. (2020)), and data augmentation (Bandara et al., 2021) have been tested to further increase the performance of GM. Furthermore, new machine learning (Godahewa et al., 2023) and deep learning (Oreshkin et al., 2020) architectures specifically designed for cross-learning have been developed. Finally, a novel area of research is emerging that tries to improve the ability of GM to capture local patterns (Sen et al., 2019), and their explainability (Rajapaksha et al., 2023).
@@
From a forecasting evaluation perspective, most of the literature on GM is focused on point prediction accuracy, possibly because most machine learning and deep learning methods do not directly output probabilistic forecasts (Makridakis et al., 2022c). Nevertheless, in many forecasting contexts (like supply chain) it is very important to be able to produce and evaluate predictions in a probabilistic way (being intervals, quantiles, or density based) (Fildes et al., 2022). Among others, Vovk et al. (2005) introduced a novel tool for uncertainty quantification under any machine learning model, namely Conformal Inference, that can also be applied in time series forecasting experiments (Stankeviciute et al., 2021).
@@
In the context of model retraining and updating strategies, instead, Spiliotis & Petropoulos (2024) is the main work related to time series forecasting. The authors explored extensively the effects of different retraining scenarios and different forms of model parameter updates on the model performance, although they focused on the exponential smoothing family of models following the traditional local approach. Huber & Stuckenschmidt (2020) briefly discussed retraining in the context of retail demand, but with few models and retraining scenarios, and on a proprietary daily dataset only. Despite the findings being promising, there has been little direct exploration of whether global models specifically require frequent retraining or if they can retain competitive accuracy with less frequent updates. However, this topic is under consideration of the broader machine learning community (Getzner et al., 2023), advocating for Green AI (Schwartz et al., 2020).
@@
Our study builds upon these existing works, directly investigating the necessity of retraining in global forecasting models. By evaluating many different retraining strategies and their impact on the forecast accuracy of several global models, we aim to provide both theoretical insights and practical recommendations for more sustainable forecasting practices.
@@ 3 Experimental design
This section describes the empirical analysis we conducted to study whether less frequent retraining scenarios may produce similar accuracy results compared to the baseline scenario (that with the highest retraining frequency). First, we describe the datasets used in the experiments, and then explain the different machine learning and deep learning models adopted. The performance measures, the several possible scenarios, and the strategy used to evaluate the forecasts are also discussed.
@@ 3.1 Datasets
For the experiment, we used two retail forecasting datasets: the M5 and the VN1 competition datasets. The M5 competition was organized by Spyros Makridakis and his colleagues as part of the M-competitions series, which aimed to compare different forecasting methods in the context of retail demand (Makridakis et al., 2022b). The M5 (Howard & Makridakis, 2020) is a well-known and well-studied dataset containing 30,490 SKU $\times$ Store daily time series of Walmart’s unit sales of products. It covers the sales of three categories of products (Food, Hobbies, and Household) sold into ten different stores located in three US states (California, Texas, and Wisconsin). The time period spans from 2011 to 2016. The time series are highly intermittent and are hierarchically organized, allowing forecasting at multiple levels, such as individual products, product categories, stores, and States. Exogenous information that can influence sales, such as product prices, promotions, and special events (e.g., holidays) are also available. The VN1 Forecasting - Accuracy Challenge competition was jointly organized by Flieber, Syrup Tech, and SupChains in October 2024, and it is the first of its series (Vandeput, 2024). The dataset contains weekly sales of 15,053 products sold from 2020 until 2024 from e-vendors. In particular, they were mostly online from the US, and the products were directly sold to the final consumers. Contrary to the M5 dataset, where all the products were sold by a single retailer (Walmart) and from just a few stores, the VN1 dataset includes product sales of 328 warehouses from 46 different retailers. As far as we know, we are among the first to test forecasting models on this data. These sets of data are the most recent and comprehensive time series datasets related to retail demand, allowing for a good generalization of the results, in particular in the context of demand forecasting.
Table 1: Characteristics of the different datasets used in the experiments. The number of series is based on the minimum number of observations per series.
@@
In both cases, in our experiment, we focused on the most disaggregated level (SKU $\times$ Store), since the potential benefit of retraining the models less frequently is much larger at lower levels of aggregation. Moreover, for the M5 competition dataset, we did not consider the whole set of time series to be able to consistently apply the evaluation setup described in Sections 3.4. In particular, for daily data, we kept only time series with at least two years of data (730 observations), while for weekly data we considered only those that had at least three years of data (157 observations).
@@ 3.2 Forecasting models
Let us define $\mathcal{Y}$ as the set of all available time series in a dataset, such that $Y_{i}$ represents a single component, and let $\mathcal{F}$ be the set of possible predictive functions, so that $F$ corresponds to a single model 444 In our setting, $F$ can be any predictive model in machine learning or deep learning framework. We do not consider classical statistical forecasting methods, like ARIMA or Exponential Smoothing, since in their common formulations are purely local models. . Without loss of generality, we can assume that all the necessary information for prediction are contained in $\mathcal{Y}$. Under the local approach, predictions for the forecast horizon $h$ are obtained training a model for each time series in the data set, implying that each time series has its own model, defined by its own parameters values.
@@
On the contrary, following the global modeling framework, forecasts for each time series are produced by a model trained on the whole data set.
@@
In our experiment, we focused on analyzing the performance of global methods only since we are interested in testing whether this approach, as opposed to the local one, can benefit from retraining the models less frequently. Indeed, nowadays, cross-learning is the go-to approach for almost all the industries involved with huge time series datasets, as in the context of retail demand forecasting, where the forecasts of many different SKUs have to be provided regularly (Januschowski et al., 2020). For a comprehensive evaluation of global modeling approaches in demand forecasting, we used both traditional machine learning models and cutting-edge deep learning methodologies. The models were selected for their established performance in time series forecasting tasks and their diverse methodological approaches, enabling a wide comparison and avoiding results being dependent on a specific algorithm. Furthermore, as benchmarking against established methods is essential in time series forecasting, we included two widely recognized global forecasting models, Linear (Pooled) Regression and Multi-Layer Perceptron, as baselines to evaluate the performance of the proposed approaches.
@@ 3.2.1 Machine learning models
Machine learning models have demonstrated their effectiveness in forecasting tasks due to their ability to capture non-linear relationships in the data. Moreover, they are often easy to train and usually produce very accurate results, leveraging the cross-learning approach. In this study, we experimented with Linear (Pooled) Regression and four different tree-based methods.
Linear Regression (LR) is a classical statistical model that assumes a linear relationship between input features and the target variable. Despite its simplicity, LR can be effective for time series forecasting when combined with appropriate feature engineering. The Pooled Regression is considered a solid benchmark for global model performance evaluation (Montero-Manso & Hyndman, 2021; Godahewa et al., 2021b), and it has also proven to be quite effective (Bandara et al., 2022).
Random Forest (RF) is an ensemble learning method based on regression trees. It constructs multiple trees during training and aggregates their predictions through averaging. RF excels in capturing non-linear patterns and interactions between variables, and it is robust to overfitting, making it a strong tree-based model for demand forecasting (Breiman, 2001). It was the method used by Amazon until 2015 to forecast e-commerce products demand (Januschowski et al., 2022) 555 For the cost of computation, the Random Forest model is tested only on the VN1 weekly dataset. .
@@
Light Gradient Boosting Machine (LGBM) is another gradient-boosted tree algorithm that emphasizes efficiency and scalability. It uses a histogram-based approach to split features and a leaf-wise growth strategy to reduce computation time. LGBM is particularly well-suited for large datasets and high-dimensional data, often outperforming other boosting algorithms in terms of speed and accuracy (Ke et al., 2017). In demand forecasting, it is among the top solutions of all the major competitions (Favorita, Rossmann, M5, and VN1) (Makridakis et al., 2022a, c; In & Jung, 2022; Lainder & Wolfinger, 2022).
Categorical Boosting (CatBoost) is a gradient-boosted decision tree algorithm specifically designed to handle categorical data. By leveraging ordered boosting and other innovations, it fosters accuracy and enhances generalization (Prokhorenkova et al., 2018). CatBoost’s ability to handle categorical features without extensive preprocessing makes it advantageous for forecasting tasks involving categorical covariates. During the last years, it consistently showed top performance on many tabular data studies, becoming the go-to solution for many practitioners in the field (Shmuel et al., 2024; Ye et al., 2025).
Machine learning models have the advantage of being easier to train with respect to their deep learning alternatives. However, they require extensive and careful feature engineering to produce accurate results (Januschowski et al., 2022). For this reason, we followed simplified versions of the M5 and VN1 top solutions to build time series features. In particular, we used lags, rolling averages, expanding averages, calendar features (year, quarter, month, week, day of week, day), and static features (store, category, location, product identifiers) based on the frequency of the dataset. Moreover, for the M5 data, we also used external features related to special events as available. The most relevant hyperparameters of the models were selected based on top-performing solutions, otherwise the default values suggested by the software providers were adopted.